# ASK: Adaptive Self-improving Knowledge Framework for Audio Text Retrieval

### v1

@@ Abstract
The dominant paradigm for Audio-Text Retrieval (ATR) relies on mini-batch-based contrastive learning. This process, however, is inherently limited by what we formalize as the Gradient Locality Bottleneck (GLB), which structurally prevents models from leveraging out-of-batch knowledge and thus impairs fine-grained and long-tail learning. While external knowledge-enhanced methods can alleviate the GLB, we identify a critical, unaddressed side effect: the Representation-Drift Mismatch (RDM), where a static knowledge base becomes progressively misaligned with the evolving model, turning guidance into noise. To address this dual challenge, we propose the Adaptive Self-improving Knowledge (ASK) framework, a model-agnostic, plug-and-play solution. ASK breaks the GLB via multi-grained knowledge injection, systematically mitigates RDM through dynamic knowledge refinement, and introduces a novel adaptive reliability weighting scheme to ensure consistent knowledge contributes to optimization. Experimental results on two benchmark datasets with superior, state-of-the-art performance justify the efficacy of our proposed ASK framework.
@@ 1 Introduction
Audio-Text Retrieval (ATR) learns a shared embedding space for audio and text (mei2022metric; yan2024bridging). The dominant paradigm relies on dual-encoder architectures trained with contrastive objectives like the NT-Xent loss (chen2020simple), which optimizes representations by exclusively contrasting samples within a mini-batch (Figure 1, left). However, the reliance on in-batch negatives is a well-recognized limitation, often failing to provide sufficiently hard negatives to effectively structure the embedding space (robinson2021can). Critically, this paradigm structurally prevents the model from leveraging any out-of-batch information, leaving the vast majority of the dataset’s semantic knowledge untapped during each optimization step.
Figure 1: Comparison between the conventional batch-only paradigm (left) and our proposed ASK framework (right).
In this work, we formalize this constraint as the Gradient Locality Bottleneck (GLB). We argue the GLB manifests in two critical failures: (1) it exacerbates semantic sparsity from under-specified text, as the model cannot access richer out-of-batch context to learn fine-grained acoustic details; and (2) it impairs long-tail generalization, a known challenge for contrastive methods (kang2020exploring), by preventing the model from forming robust decision boundaries for rare events.
A promising remedy is to augment training with an external knowledge base to access out-of-batch information (khandelwal2019generalization; guu2020retrieval). However, this introduces a critical, unaddressed challenge: a Representation-Drift Mismatch (RDM) arises as the model’s encoders evolve while the knowledge base remains static. The retrieved knowledge degrades from a source of semantic guidance to one of representational noise, destabilizing training and necessitating a co-evolution of the model and its knowledge.
@@
To systematically address this dual challenge, we propose the Adaptive Self-improving Knowledge (ASK) framework, a model-agnostic, plug-and-play solution (Figure 1, right). The ASK framework breaks the GLB by injecting information from a multi-grained knowledge base. To ensure the quality of this injection, a novel adaptive reliability weighting scheme modulates the final loss based on the cross-modal consistency of retrieved neighborhoods. Crucially, to prevent the knowledge from becoming stale, a dynamic refinement mechanism periodically updates the base, systemically mitigating RDM.
This synergistic design of reliability-governed injection and dynamic refinement proves highly effective. Extensive experiments show that ASK consistently and significantly outperforms strong baselines across multiple datasets, architectures, and interaction strategies, achieving new state-of-the-art performance.
Our main contributions are:
  • We are the first to formally define the Gradient Locality Bottleneck (GLB) in contrastive learning and the consequent Representation-Drift Mismatch (RDM) in knowledge-enhanced methods, providing rigorous mathematical formalizations for both.
  • We propose the ASK framework, a systematic solution to these challenges, featuring novel mechanisms for multi-grained knowledge injection, adaptive reliability weighting, and dynamic knowledge refinement.
@@
  • We demonstrate through extensive experiments that ASK achieves consistent state-of-the-art performance across diverse architectures and datasets, and validate the necessity of each component via comprehensive ablation studies.
@@ 2.1 Feature Representations
Feature representation serves as the cornerstone of audio-text retrieval. Early Audio-Text Retrieval (ATR) systems relied on pairing handcrafted acoustic features like MFCCs (MFCCs) with static word embeddings such as Word2Vec (Word2Vec). The advent of deep learning has led to the adoption of powerful, pre-trained unimodal encoders. Text representations are now predominantly extracted from large language models like BERT (bert), while audio features are derived from deep models pre-trained on large-scale audio datasets, such as PANNs (panns) and AST (ast). More recently, the field has shifted towards large-scale cross-modal pre-training. Models like CLAP (clap; audioclip) leverage contrastive learning on vast audio-text datasets to directly learn a shared embedding space, significantly enhancing zero-shot capabilities. Our work builds upon these advanced encoders, proposing a novel mechanism to further enhance their representations during downstream fine-tuning.
@@ 2.2 Cross-Modal Interaction and Alignment
Cross-modal interaction is key to achieving semantic alignment in ATR. Early and prevalent approaches perform this at a global, sentence-level, using contrastive learning to align the final embeddings of entire audio clips and text descriptions (clip; wav2clip; mei2022metric; liang25e_interspeech). To capture more fine-grained relationships, recent works have focused on local, token-level interactions. These methods typically employ attention mechanisms or cross-modal Transformers to model correspondences between audio frames and text tokens (scan; lu2019vilbert; xie2024gpa; yin2025atri; luo2025supclap). Our ASK framework is orthogonal to these design choices; it operates on the representations themselves and can be seamlessly integrated with both global and local interaction architectures, as demonstrated in our experiments.
@@ 3.1 Preliminaries
In a standard Audio-Text Retrieval framework, a dual-encoder architecture, comprising an audio encoder $f_{\theta}(\cdot)$ and a text encoder $g_{\phi}(\cdot)$, maps an audio-text pair $(a_{i},t_{i})$ to L2-normalized embeddings $u_{i}$ and $v_{i}$. The encoders are optimized via a symmetric NT-Xent loss (chen2020simple) over a mini-batch $B$. For a single view, the loss is:
$\mathcal{L}_{i}=-\log\frac{\exp(u_{i}^{\top}v_{i}/\tau)}{\sum_{v_{j}\in B}\exp(u_{i}^{\top}v_{j}/\tau)}$ (1)
@@ 3.2 The Gradient Locality Bottleneck
The batch-centric nature of Eq. 1 creates a fundamental limitation. To formalize this, we define the Out-of-Batch Influence (OBI) as the expected gradient norm of the loss $\mathcal{L}_{B}$ with respect to all out-of-batch embeddings:
$\text{OBI}(\mathcal{L}_{B})=\mathbb{E}_{k\in D\setminus B}\left[\left\|\frac{\partial\mathcal{L}_{B}}{\partial u_{k}}\right\|_{2}+\left\|\frac{\partial\mathcal{L}_{B}}{\partial v_{k}}\right\|_{2}\right]$ (2)
A training paradigm suffers from a Gradient Locality Bottleneck (GLB) if its OBI is identically zero, indicating no gradient flow from out-of-batch data.
For the standard contrastive loss, $\mathcal{L}_{B}$ is exclusively a function of in-batch embeddings $\{u_{j},v_{j}\}_{j\in B}$. Therefore, the partial derivatives with respect to any out-of-batch embedding $u_{k}$ or $v_{k}$ (where $k\notin B$) are necessarily zero. This directly results in $\text{OBI}(\mathcal{L}_{B})=0$, proving that standard ATR is strictly constrained by the GLB and cannot leverage the vast semantic information present in out-of-batch data.
@@ 3.3 The Representation Drift Mismatch
A direct approach to break the GLB is to perform knowledge injection, where out-of-batch knowledge is retrieved and fused with the current samples. This, however, introduces a critical challenge if the knowledge base remains static. A Representation Drift Mismatch (RDM) arises as the model’s encoders at step $t$ evolve away from the parameters used to build the base at step $t_{k}$.
To formalize this, we define RDM as the KL divergence (kullback1951information) between the ideal neighborhood distribution $P_{\text{ideal}}$ and the actual distribution $P_{\text{actual}}$. The ideal distribution is computed over a hypothetically up-to-date knowledge base, while the actual distribution uses the stale one:
$\begin{split}P_{\text{ideal}}(j|i)&=\text{softmax}_{j}(\text{sim}(f_{\theta_{t}}(a_{i}),f_{\theta_{t}}(a_{j})))\\ P_{\text{actual}}(j|i)&=\text{softmax}_{j}(\text{sim}(f_{\theta_{t}}(a_{i}),f_{\theta_{t_{k}}}(a_{j})))\end{split}$ (3)
The total RDM is then the expectation of this divergence over the dataset:
@@
$\begin{split}\text{RDM}(t,t_{k})=\mathbb{E}_{a_{i}\in D}\left[D_{KL}\left(P_{\text{ideal}}(\cdot|i)\,||\,P_{\text{actual}}(\cdot|i)\right)\right]\end{split}$ (4)
As the time difference $\Delta t=t-t_{k}$ grows, RDM increases. This corrupts the training gradients by causing a deviation in the fused knowledge vectors, $\Delta\mathcal{K}=\mathcal{K}_{\text{actual}}-\mathcal{K}_{\text{ideal}}$, where each knowledge vector $\mathcal{K}$ is the average of the Top-K retrieved embeddings.
A larger deviation in the knowledge vector $\Delta\mathcal{K}$ directly translates to a greater potential deviation in the final parameter gradients. We provide a formal proof of this entire causal chain in Appendix A. The derivation first establishes the link between the knowledge deviation $\Delta\mathcal{K}$ and the gradient deviation, and then leverages Pinsker’s inequality (cover1999elements) to bound $\|\Delta\mathcal{K}\|_{2}$ with the RDM, establishing the key relationship:
$\|\Delta\mathcal{K}\|_{2}\leq C\sqrt{2\cdot\text{RDM}(t,t_{k})}$ (5)
where $C$ is a bounded constant. Eq. 5 proves that a higher RDM widens the potential error margin for the gradient, establishing a formal link to training instability and motivating our dynamic refinement mechanism.

### v2

@@ Abstract
The dominant paradigm for Audio-Text Retrieval (ATR) relies on dual-encoder architectures optimized via mini-batch contrastive learning. However, restricting optimization to local in-batch samples creates a fundamental limitation we term the Gradient Locality Bottleneck (GLB), which prevents the resolution of acoustic ambiguities and hinders the learning of rare long-tail concepts. While external knowledge injection can break this bottleneck, it often triggers a problem called Representation-Drift Mismatch (RDM), where a static knowledge base becomes misaligned with evolving encoders, degrading guidance into noise. To address these intertwined challenges, we propose the Adaptive Self-improving Knowledge (ASK) framework. ASK breaks the GLB via multi-grained knowledge injection and mitigates RDM through a dynamic refinement strategy that synchronizes the knowledge base with the model. Additionally, an adaptive reliability weighting scheme is employed to filter retrieval noise based on cross-modal consistency. Extensive experiments across multiple benchmarks demonstrate that ASK consistently achieves new state-of-the-art performance across various backbones.
@@ 1 Introduction
Audio-Text Retrieval (ATR) aims to establish a shared embedding space where acoustically and semantically corresponding audio and text pairs are aligned (Mei et al., 2022; Yan et al., 2024). The dominant paradigm currently relies on dual-encoder architectures optimized via contrastive learning objectives, such as the NT-Xent loss (Chen et al., 2020). As illustrated in Figure 1 (left), this approach refines representations by maximizing the similarity between matched pairs while contrasting them against other samples within the same mini-batch. While effective, this in-batch mechanism implicitly restricts the optimization landscape to the local context.
The limitations of such batch-constrained optimization have been extensively discussed in broader representation learning literature, particularly regarding the necessity of large negative pools (He et al., 2020; Xiong et al., ). Building on these observations, we define this phenomenon in the context of ATR as the Gradient Locality Bottleneck (GLB). This term characterizes the specific difficulty of resolving fine-grained acoustic details that are often semantically sparse or ambiguous when the model is confined to a limited sample space. Lacking access to a broader global semantic context, the model struggles to form robust decision boundaries for long-tail events, leaving a significant portion of the dataset’s semantic potential untapped.
To break this bottleneck, one effective approach is to augment training with an external knowledge base (Dwibedi et al., 2021; Khandelwal et al., 2019; Guu et al., 2020). However, utilizing external memory introduces synchronization latency between the evolving model and the stored representations (Xiong et al., ).
Figure 1: Comparison between typical contrastive learning (left) and our proposed ASK framework (right) with a periodically updated knowledge base and an adaptive loss modulation module.
We formalize this temporal discrepancy in our continuous training framework as the Representation-Drift Mismatch (RDM). RDM describes the inevitable lag where audio and text encoders update rapidly while the external knowledge base remains static or updates slowly. Consequently, the retrieved knowledge risks degrading from a source of semantic guidance into representational noise, potentially destabilizing the training process.
To systematically address these formally defined challenges, we propose the Adaptive Self-improving Knowledge (ASK) framework. Uniquely designed as a model-agnostic enhancement, ASK can be seamlessly integrated into various backbones. As shown in Figure 1 (right), it serves as a holistic solution to inject information from a dynamically maintained multi-grained knowledge base. To handle the inherent noise and ambiguity described by the GLB, we introduce an adaptive reliability mechanism that modulates the learning signal based on cross-modal consistency. Simultaneously, to resolve the RDM, ASK employs a dynamic refinement strategy that periodically updates the knowledge base, ensuring that the external guidance co-evolves with the model.
This synergistic design enables ASK to effectively leverage global semantic structures while maintaining training stability, independent of the specific underlying architecture. Extensive experiments across multiple datasets and backbone models demonstrate that our approach consistently achieves state-of-the-art performance, validating the universality and effectiveness of our framework in handling global optimization challenges.
Our main contributions are summarized as follows:
  • We identify and define two fundamental challenges in Audio-Text Retrieval: the Gradient Locality Bottleneck (GLB), and the Representation-Drift Mismatch (RDM).
@@
  • We propose the Adaptive Self-improving Knowledge (ASK) framework, a model-agnostic solution. It breaks the GLB via multi-grained knowledge injection and systematically mitigates RDM through dynamic knowledge refinement. Furthermore, we introduce a novel adaptive reliability weighting scheme to explicitly filter retrieval noise based on cross-modal consistency.
@@
  • Extensive experiments across multiple benchmarks demonstrate that ASK consistently achieves new state-of-the-art performance across diverse global and local interaction architectures. Comprehensive ablation and zero-shot studies further validate the robustness of our framework.
@@ 2.1 Feature Representations
Feature representation serves as the cornerstone of audio-text retrieval. Early Audio-Text Retrieval systems relied on pairing handcrafted acoustic features like MFCCs (Huizen and Kurniati, 2021) with static word embeddings such as Word2Vec (Mikolov et al., 2013). The advent of deep learning has led to the adoption of powerful, pre-trained unimodal encoders. Text representations are now predominantly extracted from large language models like BERT (Devlin et al., 2019), while audio features are derived from deep models pre-trained on large-scale audio datasets, such as PANNs (Kong et al., 2020) and AST (Gong et al., 2021). More recently, the field has shifted towards large-scale cross-modal pre-training. Models like CLAP (Zhao et al., 2023; Guzhov et al., 2022) leverage contrastive learning on vast audio-text datasets to directly learn a shared embedding space, significantly enhancing zero-shot capabilities. Our work builds upon these advanced encoders, proposing a novel mechanism to further enhance their representations during downstream fine-tuning.
@@ 2.2 Cross-Modal Interaction and Alignment
Cross-modal interaction is key to achieving semantic alignment in ATR. Early and prevalent approaches perform this at a global, sentence-level, using contrastive learning to align the final embeddings of entire audio clips and text descriptions (Radford et al., 2021; Wu et al., 2022; Mei et al., 2022). To capture more fine-grained relationships, recent works have focused on local, token-level interactions. These methods typically employ attention mechanisms or cross-modal Transformers to model correspondences between audio frames and text tokens (Lee et al., 2018; Lu et al., 2019; Xie et al., 2024; Yin et al., 2025). Our ASK framework is orthogonal to these design choices; it operates on the representations themselves and can be seamlessly integrated with both global and local interaction architectures.
2.3 Retrieval-Augmented Contrastive Learning
The challenges formalized as GLB and RDM parallel fundamental bottlenecks in computer vision and information retrieval. To overcome limited in-batch negatives, frameworks like MoCo He et al. (2020) and RocketQA Qu et al. (2021) expand the contrastive denominator via memory queues, while NNCLR Dwibedi et al. (2021) utilizes external support sets. Furthermore, ANCE Xiong et al. proposes asynchronous updates to mitigate stale indices. However, these classical paradigms rely on unimodal queues or strict hard negative mining, making them highly susceptible to acoustic confusion in ATR. Unlike them, ASK structurally breaks the GLB by directly injecting out-of-batch knowledge into representations. Furthermore, to combat the subsequent representation drift, ASK introduces an adaptive reliability weighting scheme that explicitly evaluates cross-modal consistency to filter out retrieval noise.
@@ 3.1 Preliminaries
In a standard Audio-Text Retrieval framework, a dual-encoder architecture, comprising an audio encoder $f_{\theta}(\cdot)$ and a text encoder $g_{\phi}(\cdot)$, maps an audio-text pair $(a_{i},t_{i})$ to L2-normalized embeddings $u_{i}$ and $v_{i}$. The encoders are optimized via a symmetric NT-Xent loss (Chen et al., 2020) over a mini-batch $B$. For a single view, the loss is:
$\mathcal{L}_{i}=-\log\frac{\exp(u_{i}^{\top}v_{i}/\tau)}{\sum_{v_{j}\in B}\exp(u_{i}^{\top}v_{j}/\tau)},$ (1)
@@ 3.2 The Gradient Locality Bottleneck
The batch-centric nature of standard contrastive objectives creates a fundamental limitation in Audio-Text Retrieval. To formalize this, we introduce the concept of Out-of-Batch Influence (OBI), which measures the gradient contribution from data points outside the current mini-batch to the optimization process. Let $\mathcal{D}$ denote the entire training dataset and $B\subset\mathcal{D}$ represent a specific mini-batch. For a batch loss $\mathcal{L}_{B}$, the OBI is defined as the expected gradient norm with respect to all out-of-batch embeddings:
$\text{OBI}(\mathcal{L}_{B})=\mathbb{E}_{k\in\mathcal{D}\setminus B}\left[\left\|\frac{\partial\mathcal{L}_{B}}{\partial u_{k}}\right\|_{2}+\left\|\frac{\partial\mathcal{L}_{B}}{\partial v_{k}}\right\|_{2}\right],$ (2)
where $u_{k}$ and $v_{k}$ are the audio and text embeddings of an out-of-batch sample $k\in\mathcal{D}\setminus B$.
We argue that a training paradigm suffers from a Gradient Locality Bottleneck (GLB) if its OBI is identically zero, indicating that no gradient flow exists from out-of-batch data to guide the current optimization step. In a standard ATR framework using the symmetric NT-Xent loss (Eq. 1), the objective $\mathcal{L}_{B}$ is formulated exclusively as a function of the in-batch embeddings $\{u_{j},v_{j}\}_{j\in B}$. Consequently, the partial derivatives with respect to any out-of-batch embedding $u_{k}$ or $v_{k}$ (where $k\notin B$) are necessarily zero:
$\forall k\in\mathcal{D}\setminus B:\quad\frac{\partial\mathcal{L}_{B}}{\partial u_{k}}=\mathbf{0},\quad\frac{\partial\mathcal{L}_{B}}{\partial v_{k}}=\mathbf{0}.$ (3)
This directly results in $\text{OBI}(\mathcal{L}_{B})=0$, proving that standard ATR optimization is strictly constrained by the GLB and cannot leverage the vast semantic information present in out-of-batch data.
This structural confinement prevents the model from leveraging the vast semantic knowledge present in the majority of the dataset. This leads to two critical failures: (1) Semantic Ambiguity and Acoustic Hallucination, where the lack of diverse out-of-batch context prevents the model from learning fine-grained acoustic distinctions between similar but semantically distinct events; and (2) Long-tail Concept Collapse, as the reliance on limited in-batch negatives hinders the formation of robust decision boundaries for rare events, causing the model to default to common concepts with shared tonal properties.
@@ 3.3 The Representation Drift Mismatch
A direct approach to break the GLB is to perform knowledge injection, where out-of-batch samples are retrieved and fused with the current samples. By employing a simple feature fusion such as $u^{\prime}_{i}=(1-\rho)u_{i}+\rho\mathcal{K}$, where $u_{i}$ is the embedding of the current sample, $\mathcal{K}$ is the retrieved out-of-batch knowledge and $\rho$ represents the injection ratio, we establish a non-zero gradient pathway from the out-of-batch data to the model parameters. This ensures that the Out-of-Batch Influence (OBI) is no longer zero, effectively breaking the structural confinement of standard contrastive learning.
While knowledge injection breaks the GLB by introducing out-of-batch samples $\mathcal{D}\setminus B$, it inherently introduces a critical challenge: Representation Drift Mismatch (RDM). Since the encoders $f_{\theta_{t}}$ and $g_{\phi_{t}}$ are non-stationary and evolve during optimization, a static knowledge base constructed at step $t_{k}$ becomes progressively misaligned with the current model state at step $t$ ($t>t_{k}$). To formalize this, taking the audio modality as an example, we define RDM as the expected Kullback-Leibler (KL) divergence Kullback and Leibler (1951) between the ideal neighborhood distribution $P_{\text{ideal}}$ and the actual distribution $P_{\text{actual}}$:
@@
$\text{RDM}(t,t_{k})$ $=\mathbb{E}_{a_{i}\in\mathcal{D}}\left[D_{\text{KL}}\left(P_{\text{ideal}}(\cdot|i)\,||\,P_{\text{actual}}(\cdot|i)\right)\right],$ (4)
$P_{\text{ideal}}(j|i)$ $\propto\exp(\text{sim}(f_{\theta_{t}}(a_{i}),f_{\theta_{t}}(a_{j}))),$
$P_{\text{actual}}(j|i)$ $\propto\exp(\text{sim}(f_{\theta_{t}}(a_{i}),f_{\theta_{t_{k}}}(a_{j}))),$
where $a_{i}$ denotes the current query sample, the distributions are normalized over all knowledge samples indexed by $j$, $\theta_{t}$ and $\theta_{t_{k}}$ represent the model parameters at the current step $t$ and the knowledge snapshot step $t_{k}$ respectively, and $\text{sim}(\cdot,\cdot)$ is the dot product similarity.
Formally, the knowledge vector $\mathcal{K}$ is computed as the expected representation over the neighborhood distribution: $\mathcal{K}=\sum_{j}P(j)z_{j}$, where $z_{j}$ denotes the embedding of the $j$-th knowledge sample. As the training progresses, RDM accumulates and corrupts the optimization objective by inducing a deviation in the fused knowledge vectors $\mathcal{K}$, $\Delta\mathcal{K}=\mathcal{K}_{\text{actual}}-\mathcal{K}_{\text{ideal}}$.
The impact of this drift on training stability can be quantified via the gradient deviation: $\Delta\nabla=\nabla_{\theta_{t}}\mathcal{L}_{\text{actual}}-\nabla_{\theta_{t}}\mathcal{L}_{\text{ideal}}$. Specifically, consider a simplified loss: $\mathcal{L}=\mathcal{L}_{\text{main}}(u_{i},u^{\prime}_{i})$ that incorporates the knowledge-enhanced representation: $u^{\prime}_{i}=(1-\rho)u_{i}+\rho\mathcal{K}$. The gradient with respect to parameters $\theta_{t}$ is: $\nabla_{\theta_{t}}\mathcal{L}=\left(\frac{\partial\mathcal{L}}{\partial u_{i}}+(1-\rho)\frac{\partial\mathcal{L}}{\partial u^{\prime}_{i}}\right)\frac{\partial u_{i}}{\partial\theta_{t}}$. By performing a first-order Taylor expansion of the loss derivative around the ideal representation, we obtain:
$\frac{\partial\mathcal{L}_{\text{actual}}}{\partial u^{\prime}_{i}}-\frac{\partial\mathcal{L}_{\text{ideal}}}{\partial u^{\prime}_{i}}\approx H_{\mathcal{L}}(u^{\prime}_{\text{ideal}})\cdot\rho\Delta\mathcal{K},$ (5)
where $H_{\mathcal{L}}$ is the Hessian matrix. This proves that the gradient misalignment $\Delta\nabla$ is approximately proportional to the knowledge deviation $\Delta\mathcal{K}$. To bound this deviation, we leverage Pinsker’s inequality to relate the KL divergence to the Total Variation Distance, yielding the relationship:
$\|\Delta\mathcal{K}\|_{2}\leq C\sqrt{2\cdot\text{RDM}(t,t_{k})},$ (6)
where $C=\max_{j}\|z_{j}\|_{2}$ is a bounded constant. Equation 6 establishes a formal link between RDM and the potential error margin for the gradients. Higher RDM directly widens this margin, causing training instability and representational noise. This theoretical foundation necessitates our dynamic refinement mechanism, which periodically resets the RDM to zero to ensure stable co-evolution between the model and its knowledge.
Figure 2: The proposed ASK framework. A multi-grained knowledge base is periodically updated to mitigate RDM. During training, knowledge is injected into samples, and a cross-modal reliability weight is computed. A final loss is optimized using both an OT-realigned similarity matrix and the reliability weight.